40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9156     Entry time: Tue Sep 24 20:43:45 2013
Author: masayuki 
Type: Summary 
Category: SUS 
Subject: optical levers centering 

I centered optical levers of ITMX,BS,ETMY. I also change the position of optical levers of ITMX, ETMY, ITMY, BS on Friday night(9/21), of ITMX, ETMY, BS on Monday night. Both are around 6:00 ~ 7:00.But centering on Monday was totally wrong, because I centered with not good IFO alignment.

The attachment is the 5 days trend of the opt lev of ITMX. First gap is alignment on Friday and Second gap is the alignment on Monday. Now I centered after  locking the FPMI.

The attachment 2 is the last 6 hours data.  The gap on 9/25 00:00 and 1:30(UTC)  is because the alignment of the cavity and the last gap is because of  centering of the optical lever.

Attachment 1: Screenshot-Untitled_Window.png  31 kB  | Hide | Hide all
Screenshot-Untitled_Window.png
Attachment 2: Screenshot-Untitled_Window-1.png  26 kB  | Hide | Hide all
Screenshot-Untitled_Window-1.png
ELOG V3.1.3-